I need your help!
Solve the system by addition:
5x-4y= -5
x-2= -7

Answer by stanbon(75887)

Solve the system by addition:
1st: 5x-4y= -5
2nd: x-2y = -7
---------------
Multiply 2nd by 2 to get:
3rd: 2x-4y=-14
Subtract 3rd from 1st to get:
4th: 3x=9
x=3
Substitute into 2nd to get:
3-2y=-7
-2y=-10
y=5
--------
Check in 1st:
5*3-4*5=-5
15-20=-5
-5=-5
----------
Cheers,
Stan H.